OpenStudy (anonymous):

how do i find the median 25,28,29,31,32,34,36,27,37,37

OpenStudy (mitu12):

you put the numbers in order first

OpenStudy (anonymous):

i think i pretty much got it. you find the middle but what if there are two numbers in the middle

OpenStudy (mitu12):

least to greatest

OpenStudy (anonymous):

i did that

OpenStudy (sweetnikkie072):

you find the average

OpenStudy (mitu12):

then you start eliminating left to right

OpenStudy (mitu12):

you have to find the average in mean

OpenStudy (anonymous):

ok so my two middle numbers are 32 and 34 now what

OpenStudy (sweetnikkie072):

32+34/2

OpenStudy (sweetnikkie072):

=33

OpenStudy (mitu12):

i got 32

OpenStudy (anonymous):

thank you

OpenStudy (mitu12):

no problem

OpenStudy (anonymous):

its 33

OpenStudy (sweetnikkie072):

man 32+34=66/2=33

OpenStudy (mitu12):

how

OpenStudy (mitu12):

are you finding the mean

OpenStudy (anonymous):

median

OpenStudy (sweetnikkie072):

for finding mean..we have to add all the values in the given data and then divide it by the # of values given in the data

OpenStudy (sweetnikkie072):

for median we arrange in ascending order and then find the middle one..if there are to numbers there we find the average of both of them and that becomes the median

OpenStudy (sweetnikkie072):

mode is the the number which has been repeated max number of times..if none there is no mode for the given data